OpenStudy (anonymous):

3. A 5.0 µC point charge is moved within an electric field and has an electric potential energy change of 10.0 J. What is the electric potential difference before and after the charge was moved?

OpenStudy (michele_laino):

here we have to apply the subsequent formula: \[\Large \Delta U = q\Delta V\]

OpenStudy (michele_laino):

where: \[\Delta U\] is the change in energy and: \[\Delta V\] is the requested potential difference

OpenStudy (michele_laino):

q is the cahrge, of course

OpenStudy (michele_laino):

charge*

OpenStudy (michele_laino):

so, we have: \[\Large \Delta V = \frac{{\Delta U}}{q} = ...Volts\]

OpenStudy (anonymous):

Okay, so all of the things needed in the equation are in the question? So, q = 5? ΔU = 10? Am I doing that right?

OpenStudy (michele_laino):

q= 5*10^(-6) Coulombs

OpenStudy (anonymous):

Oh, okay. So then ΔV = 10? Or..

OpenStudy (michele_laino):

\[\Large \Delta V = \frac{{\Delta U}}{q} = \frac{{10}}{{5 \times {{10}^{ - 6}}}} = \frac{{10}}{5} \times {10^6} = ...Volts\]

OpenStudy (michele_laino):

@jaysabelle

OpenStudy (anonymous):

So it would be \[2x10^{6}\] = volts?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

your answer is correct!

OpenStudy (anonymous):

Really? Thank you!